OpenStudy (anonymous):

Check my work: Find the volume of the solid obtained by rotating the region bounded by the curves y = x and y = x^2 about the y-axis.

OpenStudy (anonymous):

y = x -> x = y y = sqrt(y) when y = 0,1 y = x^2 -> x = sqrt(y) \[\pi \int\limits_{0}^{1} (\sqrt{y})^2 - (y)^2 = \pi/6\]

OpenStudy (anonymous):

@experimentX @hartnn

OpenStudy (experimentx):

I think it's correct.

OpenStudy (experimentx):

i mean the procedure.

OpenStudy (anonymous):

what I meant to write was integral (sqrty) -(y)^2 from 0 to 1. Is that good?

OpenStudy (anonymous):

@experimentX